Ram’s father is thrice as old as Ram was 2 years ago. Five years from now, his father’s age will be 6 years more than twice the ram’s age. What is Ram’s present age?
48yrs
40yrs
45yrs
None

Answer by lwsshak3(11628)

Ram’s father is thrice as old as Ram was 2 years ago. Five years from now, his father’s age will be 6 years more than twice the ram’s age. What is Ram’s present age?
48yrs
40yrs
45yrs
None
**
let x=Ram's present age
2 yrs ago:
x-2 was Ram's age
3(x-2)=Father's present age
..
5 yrs from now:
3(x-2)+5=father's age
x+5=ram's age
3(x-2)+5=2(x+5)+6
3x-6+5=2x+10+6
x=17
3(x-2)=3*15=45
Ram's present age=17
Father's present age=45